OpenStudy (anonymous):

How many solutions does the system of equations have? x = 6y + 6 2x – 12y = 12

OpenStudy (luigi0210):

If you the 12 over, and divide by 2, you will see that the equations are the same. Meaning they have infinitely many soultions~

OpenStudy (luigi0210):

*solutions

OpenStudy (ajanijones):

I know the answer.....

OpenStudy (anonymous):

@ajanijones what is it ?

OpenStudy (anonymous):

thank youuu @Luigi0210 thank youuu

OpenStudy (anonymous):

8) How many solutions does the system of equations have? 2x + 3y = 12 y – 3x = 1

OpenStudy (ajanijones):

I no tell bye.

OpenStudy (anonymous):

9) Solve the system using substitution. x = 2y + 6 x + y = 12

OpenStudy (anonymous):

@Luigi0210 can you help with these other ones

OpenStudy (anonymous):

Ur answer is Listen to @Luigi0210